OpenStudy (anonymous):

6. Is the line with equation y = -8 horizontal or vertical? What is the slope of this line?

OpenStudy (saifoo.khan):

totally horizontal.

OpenStudy (saifoo.khan):



OpenStudy (across):

No, it's a curve defined in\[\mathbb{R}^{16}\]that can be expressed using elliptic equations.

OpenStudy (across):

... my point is:

OpenStudy (across):

andres123, a line is defined as follows:\[y=mx+b.\]I pray you've seen that equation before. Now, if there is no x in y = mx + b, what does that imply? It implies m = 0. What does a slope (m) of 0 indicate?

OpenStudy (saifoo.khan):

\[\huge \color\white{i -guess- this- is -your- point???}\]

OpenStudy (across):

m is taught as being "rise over run." When is a fraction 0?\[\frac{0}{c},\] where c is any number. Thus, if the "rise" is 0, then the line is horizontal. I hope you got that.

OpenStudy (anonymous):

no.... acrosss.... your so lost... not even a map will help you right now... its ok though

OpenStudy (anonymous):

Actually, across is exactly right and is trying to show you something important here.